OpenStudy (anonymous):

****************medals and fans*****************Samantha donated 10 T-shirts to charity and has 16 left. Which equation could be used to find x, the number of T-shirts Samantha had originally?

OpenStudy (anonymous):

x - 10 = -16 x + 10 = 16 x + 16 = -10 x - 16 = 10

OpenStudy (anonymous):

x-16=10

OpenStudy (anonymous):

Thank you I know my husband will enjoy this answer haaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa

OpenStudy (anonymous):

is that all

OpenStudy (anonymous):

yes @undockedraptor

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

bye haaaaaaaaaaaaaaaaaaaaaaa